1.     The following production function has been estimated for wheat:


where, W = Wheat yield (kg/ha), F = Fertiliser (kg/ha) and L = Labour (hours/ha)

Derive an expression for the least-cost expansion path given that fertiliser costs R12 per kg and labour costs R6 per hour.                                                                                       


1
Expert's answer
2021-09-03T08:09:34-0400

The complete question is:

 The following production function has been estimated for wheat:

"w=4F^{0.5}.L^{0.25}"

where, W = Wheat yield (kg/ha), F = Fertiliser (kg/ha) and L = Labour (hours/ha)

Derive an expression for the least-cost expansion path given that fertiliser costs R12 per kg and labour costs R6 per hour.       


Solution


At equilibrium, MRTS "= \frac{Price \space of\space fertilizer}{Price \space of\space Labor}"

"MRTS = \frac{MPF }{ MPL}"


"MPF = \frac{dW}{dF} = 4\times0.5(\frac{L^{0.25}}{F^{0.5}}) = 2\times\frac{L^{0.25}}{F^{0.5}}"


"MPL = \frac{dW}{dL} = 4\times0.25\frac{F^{0.5}}{L^{0.75}} = \frac{F^{0.5}}{L^{0.75}}"


"MRTS = 2\times\frac{L^{0.25}}{F^{0.5}} \times\frac{F^{0.5}}{L^{0.75}}"


"MRTS = 2(\frac{L}{F})"


MRTS "= \frac{Price \space of\space fertilizer}{Price \space of\space Labor}"


"2(\frac{L}{F}) = \frac{12 }{ 6} = 2"

L = F : equation of expansion path